# Non-Expansive Mappings in Two-Time-Scale Stochastic Approximation: Finite-Time Analysis

### v1

@@ 4 Gradient Descent Operators in Slower Time-Scale
In this case, the function $g(x^{*}(\cdot),\cdot)$ can be shown to be contractive, and hence we have a two-time-scale iteration where both time-scales have contractive mappings. In this case, the result from [10] can directly be applied, to obtain a bound of $\mathcal{O}(1/k^{2/3})$ on $\mathbb{E}[\|x_{k}-x^{*}(y_{k})\|^{2}+\|y_{k}-y^{*}\|^{2}]$. Here $y^{*}$ is the unique minimizer for $J(\cdot)$.
@@
    The stepsize sequences satisfy $\mathfrak{a}<\mathfrak{b}<1$ and $0.5<\mathfrak{b}$. We also have the key assumption that $\beta_{k}^{2}\leq\alpha_{k}^{3}$ and $\beta_{k}\alpha_{k}\leq\beta_{0}\alpha_{0}/(k+1)$.
The key difference in assumptions is that we no longer require $\alpha_{k}$ to be square summable, allowing sequences such as $\alpha_{k}=\mathcal{O}(1/k^{0.4})$ and $\beta_{k}=\mathcal{O}(1/k^{0.6})$. Then, we have the following theorem, with its proof presented in Appendix C.
@@
Suppose that Assumptions 4 and 4 hold. Then there exist constants $C_{3},C_{4}$ such that if $\alpha_{0}\leq C_{4}$, then for all $k\geq 0$,
@@
The optimal rate in this case is obtained at $\mathfrak{a}=0.4$ and $\mathfrak{b}=0.6$, with a bound of $\min_{0\leq i\leq k}\mathbb{E}\left[\|\nabla J(y_{i})\|^{2}\right]=\mathcal{O}(\log(k)/k^{0.4})$. In this case, we do not have a bound on the gradient at time $k$, but only have a bound on the minimum gradient attained before time $k$. Such non-convex two-time-scale algorithms arise in non-convex minimax optimization and reinforcement learning (actor-critic algorithms) with similar bounds obtained in [29, 28] under different settings.
@@ 5.1 Minimax Optimization
$\min_{y\in\mathbb{R}^{d_{2}}}\max_{x\in\mathbb{R}^{d_{1}}}H(x,y),$
where $H:\mathbb{R}^{d_{1}}\times\mathbb{R}^{d_{2}}\mapsto\mathbb{R}$ is the objective function. This formulation of finding saddle point solutions to minimax optimization problems arises in several areas, including two-player zero-sum games, generative adversarial networks, and robust statistics [20]. We consider the Two-Time-scale Stochastic Gradient Descent Ascent (TTSGDA) algorithm [20] for finding saddle points.
(3) $x_{k+1}=x_{k}+\alpha_{k}(\nabla_{x}H(x_{k},y_{k})+M_{k+1})$
@@
We work under the setting of smooth strongly concave-convex minimax optimization. This is formally stated in the following assumption. {assumption} Function $H(x,y)$ is $L-$smooth, $\rho-$strongly concave in $x$ and convex in $y$. This assumption is sufficient to ensure that the faster time-scale (update of iterate $x_{k}$) has a contractive mapping and the slower time-scale (update of iterate $y_{k}$) has a non-expansive mapping. Define $\Phi(y)=\max_{x\in\mathbb{R}^{d_{1}}}H(x,y)$. Then we have the following result, with its proof presented in Appendix D.
@@
Suppose Assumption 5.1 holds for iterates $\{x_{k},y_{k}\}$ defined in (3), and let $\alpha_{k}=1/(k+1)^{0.5+0.66\epsilon}$ and $\beta_{k}=1/(k+1)^{0.75+\epsilon}$. Then the iterates $\{x_{k},y_{k}\}$ converge to the set of saddle points of the function $H(x,y)$ a.s., and $\mathbb{E}\left[\|\nabla\Phi(y)\|^{2}\right]\leq\mathcal{O}(1/k^{1/4-\epsilon}).$ Here $\epsilon$ can be arbitrarily small.
We also want to mention that a similar result can be obtained for the case where $x_{k}$ is bounded in a convex and bounded set $\mathcal{X}$ using the framework from Subsection 2.1.
@@ 5.2 Linear Stochastic Approximation
This algorithm finds use in reinforcement learning [16] and linear SA with Polyak averaging [18]. Define matrix $\Delta=A_{22}-A_{21}A_{11}^{-1}A_{12}$. Under the assumption that matrices $-A_{11}$ and $-\Delta$ are Hurwitz, i.e., the real part of all their eigenvalues is negative, iteration (5) can be written as two-time-scale SA with both time-scales as fixed point iterations with contractive mappings.
In this paper, we consider the case where the matrix $-A_{11}$ is negative definite and the matrix $-\Delta$ is negative semidefinite. Formally, we make the following assumption. {assumption} Iteration (5) satisfies the following.
@@
    The matrix $-A_{11}$ is negative definite and the matrix $-\Delta$ is (non-zero) negative semidefinite.
@@
Part a) of this assumption ensures that iteration (5) satisfies Assumptions 2 and 2. As Hurwitz matrices also correspond to contractive fixed point iterations [18], we can also work with Hurwitz $-A_{11}$ (instead of negative definite). Here is the result for this setting, with its proof presented in Appendix D.
@@
Suppose Assumption 5.2 holds for iterates $\{x_{k},y_{k}\}$ defined in (5), and let $\alpha_{k}=1/(k+1)^{0.5+0.66\epsilon}$ and $\beta_{k}=1/(k+1)^{0.75+\epsilon}$. Then the iterates $\{x_{k},y_{k}\}$ converge to the set of solutions of system of equations (4) a.s., and
@@ 5.3 Constrained Optimization with Lagrangian Multipliers
This algorithm is particularly useful in distributed settings, where each node or agent updates some part of the variable $x$ under some local convex constraints, and there is a global linear constraint. In that case, the local updates only require knowledge of the gradient, the Lagrange multiplier, and their local constraints. Another similar setting in which a similar algorithm can be used is generalized Nash equilibrium problems (or Nash equilibrium under constraints) [14] and game control for strongly monotone games [5]. In this setting, each player takes action in a convex set and updates their action using gradient ascent. The game controller has a global linear objective which they want the player’s actions to satisfy [5].
@@
    $-Ax^{*}(\lambda)$ is strongly monotone.
@@
Suppose Assumption 5.3 holds for iterates $\{x_{k},\lambda_{k}\}$ defined in (7), and let $\alpha_{k}=1/(k+1)^{0.5+0.66\epsilon}$ and $\beta_{k}=1/(k+1)^{0.75+\epsilon}$. Then $x_{k}$ converges to $\hat{x}(\lambda_{k})$ and $Ax_{k}$ converges to $b_{0}$ a.s., and
@@ 6 Conclusion
In this paper, we studied two-time-scale iterations where the slower time-scale has a non-expansive mapping. Our main contribution is the finite-time bound for this setting. We showed that the mean square error decays at a rate of $\mathcal{O}(k^{-1/4+\epsilon})$, where $\epsilon>0$ is arbitrarily small depending on the step size. Our analysis takes inspiration from techniques in contractive SA and inexact KM iterations. Extensions of our work include studying the same algorithm under Markovian noise and extending the result to Banach spaces. Another extension could be to study non-expansive mappings in both time-scales, but the analysis would require further assumptions to ensure that the faster time-scale has a unique fixed point for each $y$.
@@
We first use Corollary 2.14 from [1] which states that for any $\epsilon\in(0,1)$ and $y_{1},y_{2}\in\mathbb{R}^{d_{2}}$, we have $\|(1-\epsilon)y_{1}+\epsilon y_{2}\|^{2}=(1-\epsilon)\|y_{1}\|^{2}+\epsilon\|y_{2}\|^{2}-\epsilon(1-\epsilon)\|y_{1}-y_{2}\|^{2}.$
@@
Here we have dropped the term $-\beta_{k}(1-\beta_{k})\mathbb{E}\left[\|y_{k}-g(x_{k},y_{k})\|^{2}\right]$ as $10<\beta_{k}<$ 1. Expanding the recursion, we get the following.
@@
We have also shown that $x^{*}(y)$ is Lipschitz in $y$ (Lemma 3.1). And hence using Lemma 8.1 from [2], $x_{t}\rightarrow x^{*}(y_{t})$ with probability $1$. Now consider the ODE limit
@@
For the first term above, there exists $D_{2}$ such that $1-\mu^{\prime}\alpha_{0}/(k+1)^{\mathfrak{a}}\leq 1-1/(k+D_{2})$. Using Corollary 2.1.2 from [7], we have
$\prod_{i=0}^{k-1}(1-\mu^{\prime}\alpha_{i})\leq\prod_{i=0}^{k-1}\left(1-\frac{1}{i+D_{2}}\right)\leq\frac{D_{2}}{k+D_{2}}.$
@@
Here we use the fact that $\alpha_{k}$ is of the form $\alpha_{0}/(k+1)^{\mathfrak{a}}$ where $10<\mathfrak{a}<$ 1. Hence we get the following bound.
(15) $\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]\leq\frac{C_{1}}{(k+1)^{\mathfrak{a}}}=D_{3}\alpha_{k}.$
Here $C_{1}=D_{3}\alpha_{0}=\kappa_{2}\left(1+\|x_{0}-x^{*}(y_{0})\|^{2}+\|y_{0}-y^{*}\|^{2}\right)$ where $\kappa_{2}=\max\{\mu^{\prime}D_{2}+2\Gamma_{1}\alpha_{0}\kappa_{1},2\Gamma_{1}\alpha_{0}(1+\kappa_{1}^{2})\}/\mu^{\prime}$. This completes our bound on $\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]$.
@@
$\mathbb{E}[\|y_{k}-z_{k}\|^{2}]\leq\mathbb{E}\left[\|U_{k}\|^{2}\right]$ $\leq\sum_{i=0}^{k-1}\left(\beta_{i}\prod_{j=i+1}^{k-1}(1-\beta_{j})\right)^{2}\mathbb{E}\left[\|M^{\prime}_{i+1}\|^{2}\right]$
@@
Here, inequality (a) follows from Assumption 2 and Lemma 3.3, which together imply that $M^{\prime}_{i+1}$ is bounded in expectation by some constant $D_{4}$. Inequality (b) follows from Corollary 2.1.2 from [7].
@@
[…]
$\stackrel{{\scriptstyle(a)}}{{\leq}}\alpha_{0}\beta_{0}\left(\frac{1}{(i+1)^{\mathfrak{a}+\mathfrak{b}}}+\frac{1}{(\mathfrak{a}+\mathfrak{b}-1)(i+1)^{\mathfrak{a}+\mathfrak{b}-1}}\right)\stackrel{{\scriptstyle(b)}}{{\leq}}\frac{2\alpha_{0}\beta_{0}}{\mathfrak{a}+\mathfrak{b}-1}\frac{1}{(i+1)^{\mathfrak{a}+\mathfrak{b}-1}}.$
Here, inequality (a) and (b) both follow from the fact that $21<\mathfrak{a}+\mathfrak{b}<$ 2, which follows from our assumption that $\alpha_{k}$ and $\beta_{k}$ are both non-summable but square summable sequences. Next,
$\sum_{i=0}^{k}\beta_{i}\sum_{j=i}^{k-1}\beta_{j}\alpha_{j}\leq\frac{2\alpha_{0}\beta_{0}^{2}}{\mathfrak{a}+\mathfrak{b}-1}\sum_{i=0}^{k-1}\frac{1}{(i+1)^{\mathfrak{a}+2\mathfrak{b}-1}}\stackrel{{\scriptstyle(c)}}{{\leq}}\frac{2\alpha_{0}\beta_{0}^{2}}{\mathfrak{a}+\mathfrak{b}-1}\sum_{i=0}^{\infty}\frac{1}{(i+1)^{4\mathfrak{a}-1}}\stackrel{{\scriptstyle(d)}}{{\eqqcolon}}D_{5}.$
Inequality (c) follows from the fact that $2\mathfrak{b}\leq 3\mathfrak{a}$, which follows from our assumption that $\beta_{k}^{2}\leq\alpha_{k}^{3}$. Finally, inequality (d) follows from the fact that $14\mathfrak{a}-1>$ 1, which follows from our assumption that $\mathfrak{a}>0.5$ or $\alpha_{k}$ is square summable.
@@
where $C_{2}\coloneqq(2\Gamma_{6}+2\Gamma_{7}D_{5})/\beta_{0}+8\Gamma_{5}\beta_{0}$. Here, the first inequality follows from (A.5), and the last inequality follows from the fact that $1/(k+1)^{\mathfrak{b}}\leq 1/(k+1)^{1-\mathfrak{b}}$, which follows from our assumption that $\mathfrak{b}>1/2$. This completes the proof.

### v3

@@ 4 Gradient Descent Operators in Slower Time-Scale
In this case, the function $g(x^{*}(\cdot),\cdot)$ can be shown to be contractive, and hence we have a two-time-scale iteration where both time-scales have contractive mappings. In this case, the result from [5] can directly be applied, to obtain a bound of $\mathcal{O}(1/k)$ on $\mathbb{E}[\|x_{k}-x^{*}(y_{k})\|^{2}+\|y_{k}-y^{*}\|^{2}]$. Here $y^{*}$ is the unique minimizer for $J(\cdot)$.
@@
    The stepsize sequences satisfy $\mathfrak{a}<\mathfrak{b}<1$ and $0.5<\mathfrak{b}$. We also have the key assumption that $\beta_{k}^{2}\leq\alpha_{k}^{3}$ and $\beta_{k}\alpha_{k}\leq\beta\alpha/(k+K_{1})$.
The key difference in assumptions is that we no longer require $\alpha_{k}$ to be square summable, allowing sequences such as $\alpha_{k}=\mathcal{O}(1/k^{0.4})$ and $\beta_{k}=\mathcal{O}(1/k^{0.6})$. Then, we have the following theorem, with its proof presented in Appendix D.
@@
Suppose that Assumptions 4 and 4 hold. Then there exist constants $C_{3},C_{4}$ such that if $\alpha\leq C_{4}$, then for all $k\geq 0$,
@@
The optimal rate in this case is obtained at $\mathfrak{a}=0.4$ and $\mathfrak{b}=0.6$, with a bound of $\min_{0\leq i\leq k}\mathbb{E}\left[\|\nabla J(y_{i})\|^{2}\right]=\mathcal{O}(\log(k)/k^{0.4})$. In this case, we do not have a bound on the gradient at time $k$, but only have a bound on the minimum gradient attained before time $k$. Such non-convex two-time-scale algorithms arise in non-convex minimax optimization and reinforcement learning (actor-critic algorithms). The same upper bound of $\mathcal{O}(\log(k)/k^{0.4})$ was obtained in [35, 33] under different assumptions on the noise sequence, and remains the best-known upper bound for this setting.
@@ 5.1 Minimax Optimization
$\min_{y\in\mathbb{R}^{d_{2}}}\max_{x\in\mathcal{X}}H(x,y),$
where $H:\mathcal{X}\times\mathbb{R}^{d_{2}}\mapsto\mathbb{R}$ is the objective function. This formulation of finding saddle point solutions to minimax optimization problems arises in several areas, including two-player zero-sum games, generative adversarial networks, and robust statistics [24]. We consider the Two-Time-scale Stochastic Gradient Descent Ascent (TTSGDA) algorithm [24] for finding saddle points.
(3) $x_{k+1}=\mathcal{P}_{\mathcal{X}}\left(x_{k}+\alpha_{k}(\nabla_{x}H(x_{k},y_{k})+M_{k+1})\right)$
@@
We work under the setting of smooth strongly concave-convex minimax optimization. This is formally stated in the following assumption. {assumption} Function $H(x,y)$ is $L$-smooth, $\rho$-strongly concave in $x$ and convex in $y$. The set $\mathcal{X}\subset\mathbb{R}^{d_{1}}$ is a convex set with a bounded diameter. This assumption is sufficient to ensure that the faster time-scale (update of iterate $x_{k}$) has a contractive mapping and the slower time-scale (update of iterate $y_{k}$) has a non-expansive mapping. Define $\Phi(y)=\max_{x\in\mathbb{R}^{d_{1}}}H(x,y)$. Then we have the following result, with its proof presented in Appendix B.
@@
Suppose Assumption 5.1 holds for iterates $\{x_{k},y_{k}\}$ defined in (3), and let $\alpha_{k}=\mathcal{O}\left(1/(k+1)^{0.5+0.66\epsilon}\right)$ and $\beta_{k}=\mathcal{O}\left(1/(k+1)^{0.75+\epsilon}\right)$. Then the iterates $\{x_{k},y_{k}\}$ converge to the set of saddle points of the function $H(x,y)$ a.s., and $\mathbb{E}\left[\|\nabla\Phi(y_{k})\|^{2}\right]\leq\mathcal{O}(1/k^{1/4-\epsilon}).$ Here $\epsilon$ can be arbitrarily small.
We note that stronger guarantees can be obtained for strongly convex-concave minimax optimization by using techniques such as extra-gradient, proximal best response, or just by considering averaged iterates [9, 34], but to the best of our knowledge, our result is the first last-iterate bound for vanilla TTSGDA.
@@ 5.2 Linear Stochastic Approximation
This algorithm finds use in reinforcement learning [20] and linear SA with Polyak averaging [22]. Define matrix $\Delta=A_{22}-A_{21}A_{11}^{-1}A_{12}$. Under the assumption that matrices $A_{11}$ and $\Delta$ are Hurwitz, i.e., the real part of all their eigenvalues is negative, iteration (5) can be written as two-time-scale SA with both time-scales as fixed point iterations with contractive mappings.
In this paper, we consider the case where the matrix $A_{11}$ is negative definite and the matrix $\Delta$ is negative semidefinite. Formally, we make the following assumption. {assumption} Iteration (5) satisfies the following.
@@
    The matrix $A_{11}$ is negative definite and the matrix $\Delta$ is (non-zero) negative semidefinite.
@@
Part a) of this assumption ensures that iteration (5) satisfies Assumptions 2 and 2. As Hurwitz matrices also correspond to contractive fixed point iterations [22], we can also work with Hurwitz $A_{11}$ (instead of negative definite). Here is the result for this setting, with its proof presented in Appendix B.
@@
Suppose Assumption 5.2 holds for iterates $\{x_{k},y_{k}\}$ defined in (5), and let $\alpha_{k}=\mathcal{O}\left(1/(k+1)^{0.5+0.66\epsilon}\right)$ and $\beta_{k}=\mathcal{O}\left(1/(k+1)^{0.75+\epsilon}\right)$. Then the iterates $\{x_{k},y_{k}\}$ converge to the set of solutions of system of equations (4) a.s., and
@@
We note that the linear non-expansive map can also be treated as a linear contractive map under the seminorm $\|x\|_{\Delta}\coloneqq\sqrt{x^{\top}\Delta x}$, which vanishes on $\ker(\Delta)$. We believe that stronger guarantees can be obtained for iteration (5) using existing analyses for seminorm-contractive stochastic approximation [12], although such analyses are currently available only for the single-time-scale setting and would need to be extended to the two-time-scale case. We include this application to show the broad applicability of our framework.
@@ 5.3 Constrained Optimization with Lagrangian Multipliers
This problem can be viewed as a special case of strongly convex–concave minimax optimization. We treat it separately due to the interesting role played by projections (Lemma 5.3), as well as its practical applications. The algorithm is particularly useful in distributed settings, where each node or agent updates some part of the variable $x$ under some local convex constraints, and there is a global linear constraint. In that case, the local updates only require knowledge of the gradient, the Lagrange multiplier, and their local constraints. Another similar setting in which a similar algorithm can be used is generalized Nash equilibrium problems (or Nash equilibrium under constraints) [18] and game control for strongly monotone games [6]. In this setting, each player takes action in a convex set and updates their action using gradient ascent. The game controller has a global linear objective which they want the player’s actions to satisfy [6].
@@
    $Ax^{*}(\lambda)$ is strongly monotone.
@@
Suppose Assumption 5.3 holds for iterates $\{x_{k},\lambda_{k}\}$ defined in (7), and let $\alpha_{k}=\mathcal{O}\left(1/(k+1)^{0.5+0.66\epsilon}\right)$ and $\beta_{k}=\mathcal{O}\left(1/(k+1)^{0.75+\epsilon}\right)$. Then $x_{k}$ converges to $\hat{x}(\lambda_{k})$ and $Ax_{k}$ converges to $b_{0}$ a.s., and
@@
6 Numerical Experiments
In this section, we simulate the three applications from the previous section. Each plot represents the average over 200 independent runs of the algorithm. The shaded region indicates one standard deviation around the mean. The stepsize sequences are of the form $\alpha_{k}=\alpha/(k+100)^{\mathfrak{a}}$ and $\beta_{k}=\beta/(k+100)^{\mathfrak{b}}$.
In Figure 1, we simulate the linear stochastic approximation setting. We generate a random positive definite matrix $A_{11}\in\mathbb{R}^{20\times 20}$ and generate a positive semi-definite matrix $\Delta\in\mathbb{R}^{20\times 20}$ of rank $5$. We set $A_{12}=A_{21}=I$ and $A_{22}=A_{11}^{-1}+\Delta$. The vectors $b_{1},b_{2}\in\mathbb{R}^{20}$ are obtained by randomly generating $x^{\prime},y^{\prime}\in\mathbb{R}^{20}$ and computing $b_{1}=A_{11}x^{\prime}+A_{12}y^{\prime}$ and $b_{2}=A_{21}x^{\prime}+A_{22}y^{\prime}$. At each timestep, i.i.d. normal noise is added to the vectors and the matrices, and the noisy residual error is used for updating the iterates $x_{k}$ and $y_{k}$.
In Figure 1(a), we study the effect of the exponent in the stepsize (parameters $\mathfrak{a}$ and $\mathfrak{b}$) on the residual error. We see that a slowly decreasing stepsize has a better rate of convergence. This matches our intuition and the bound obtained in Theorem 2.1. In Figure 1(b), we study the effect of constants $\alpha,\beta$ on the residual error. As before, larger stepsizes accelerate convergence. Interestingly, the residual error appears to be primarily influenced by the value of $\beta$, with minimal sensitivity to $\alpha$.
(a) Variation with exponents $\mathfrak{a},\mathfrak{b}$
(b) Variation with constants $\alpha,\beta$
Figure 1: Linear Stochastic Approximation: effect of stepsize sequence on residual errors
We also want to note an interesting pattern observed consistently across all three scenarios. Even though the choices $\mathfrak{a}=0.5,\mathfrak{b}=0.5$ and $\mathfrak{a}=0.4,\mathfrak{b}=0.6$ do not satisfy the assumptions required for our theoretical bounds, they show the the fastest empirical convergence in Figures 13. This may indicate that stronger bounds may be possible under modified assumptions. However, these settings also exhibit sensitivity to parameter perturbations: for instance, increasing $\alpha$ from $0.1$ to $0.5$ can lead to instability in some runs, with iterates diverging to magnitudes exceeding $10^{200}$.
(a) Residual error for the faster time-scale
(b) Residual error for the slower time-scale
Figure 2: Minimax Optimization: residual errors for the two time-scales
We present the simulations for minimax optimization in Figure 2. We choose the function $H(x,y)=x^{T}Ay-0.5\|x\|^{2}+(y^{T}Qy)^{2}$ where $A\in\mathbb{R}^{5\times 5}$ is a randomly generated matrix and $Q\in\mathbb{R}^{5\times 5}$ is a randomly generated symmetric positive semi-definite matrix. In Figure 2(a), we plot $\|\nabla_{x}H(x_{k},y_{k})\|$, the residual error for the faster time-scale and in Figure 2(b), we plot $\|\nabla\Phi(y_{k})\|$, the residual error for the slower time-scale. We observe that both these residual errors converge to zero, and the dependence on the stepsize is similar to that observed in the linear case.
For constrained optimization, we choose the function $H_{0}(x)=\|x-\ell\|^{2}+\sum_{i=1}^{20}e^{x(i)}$ where $x,\ell\in\mathbb{R}^{20}$ and $x(i)$ denotes the $i$-th element of the vector. Motivated by the application to distributed optimization, we impose local constraints of the form $H_{j}(x)\leq 0$ for each $j\in{1,\ldots,4}$, where $H_{j}(x)=\sum_{i=5j+1}^{5j+5}x(i)^{2}-4$. These constraints enforce that each 5-dimensional block of the vector $x$ lies within a ball of radius two. Vectors $b_{0},\ell$ and the matrix $A$ are chosen randomly to ensure that all assumptions are satisfied. In Figure 3(a), we study the effect of the stepsize constants $\alpha$ and $\beta$ on the residual error. Since the faster time-scale is projected, larger stepsizes can be used without causing instability in the iterates. However, while increasing the stepsize initially improves convergence rate, further increases eventually degrade performance.
(a) Effect of constants $\alpha,\beta$
(b) Comparison with the non-projected case
Figure 3: Constrained optimization: (a) effect of stepsize on residual error, and (b) comparison with the case where there is no projection in the faster time-scale
In Figure 3(b), we compare the performance of the projected case with the case where there are no projections in the faster time-scale. Recall that the projection in the faster time-scale leads to the non-expansiveness in the slower time-scale and in the absence of a projection, both time-scales have contractive mappings (Lemma 5.3). We observe that the no-projection case exhibits faster convergence, aligning with the intuition that contractive mappings lead to improved convergence rates (Table 1).
@@ 7 Conclusion
In this paper, we studied two-time-scale iterations where the slower time-scale has a non-expansive mapping. Our main contribution is the finite-time bound for this setting. We showed that the mean square residual error decays at a rate of $\mathcal{O}(k^{-1/4+\epsilon})$, where $\epsilon>0$ is arbitrarily small depending on the step size. Our analysis takes inspiration from techniques in contractive SA and inexact KM iterations. Extensions of our work include studying the same algorithm under Markovian noise and extending the result to Banach spaces. Another extension could be to study non-expansive mappings in both time-scales, but the analysis would require further assumptions to ensure that the faster time-scale has a unique fixed point for each $y$.
@@
We first use Corollary 2.14 from [1] which states that for any $\epsilon\in(0,1)$ and $y_{1},y_{2}\in\mathbb{R}^{d_{2}}$, we have $\|(1-\epsilon)y_{1}+\epsilon y_{2}\|^{2}=(1-\epsilon)\|y_{1}\|^{2}+\epsilon\|y_{2}\|^{2}-\epsilon(1-\epsilon)\|y_{1}-y_{2}\|^{2}.$ Therefore,
@@
Here we have dropped the term $\beta_{k}(1-\beta_{k})\mathbb{E}\left[\|y_{k}-g(x_{k},y_{k})\|^{2}\right]$ as $0<\beta_{k}<1$. Expanding the recursion, we get the following.
@@
We have also shown that $x^{*}(y)$ is Lipschitz in $y$ (Lemma 3.1). And hence using Lemma 8.1 from [2], $\|x_{t}-x^{*}(y_{t})\|\rightarrow 0$ with probability $1$. Now consider the ODE limit
@@
For the first term above, there exists $D_{2}$ such that $1-\mu^{\prime}\alpha/(k+K_{1})^{\mathfrak{a}}\leq 1-1/(k+1+D_{2})$. Using Corollary 2.1.2 from [10], we have
$\prod_{i=0}^{k-1}(1-\mu^{\prime}\alpha_{i})\leq\prod_{i=0}^{k-1}\left(1-\frac{1}{i+1+D_{2}}\right)\leq\frac{D_{2}}{k+D_{2}}.$
@@
Here we use the fact that $\alpha_{k}$ is of the form $\alpha/(k+K_{1})^{\mathfrak{a}}$ where $0<\mathfrak{a}<1$ and $K_{1}\geq\left(\frac{2\mathfrak{a}}{\mu\alpha}\right)^{1-\mathfrak{a}}$. Hence we get the following bound.
(15) $\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]\leq\frac{C_{1}}{(k+K_{1})^{\mathfrak{a}}}=D_{3}\alpha_{k}\leq\frac{C_{1}}{(k+1)^{\mathfrak{a}}}.$
Here $C_{1}=D_{3}\alpha=\kappa_{2}\left(1+\|x_{0}-x^{*}(y_{0})\|^{2}+\|y_{0}-y^{*}\|^{2}\right)$ where $\kappa_{2}=\max\{\mu^{\prime}D_{2}+2\Gamma_{1}\alpha\kappa_{1},2\Gamma_{1}\alpha(1+\kappa_{1}^{2})\}/\mu^{\prime}$. This completes our bound on $\mathbb{E}\left[\|x_{k}-x^{*}(y_{k})\|^{2}\right]$.
@@
$\mathbb{E}[\|y_{k}-z_{k}\|^{2}]=\mathbb{E}\left[\|U_{k}\|^{2}\right]$ $\leq\sum_{i=0}^{k-1}\left(\beta_{i}\prod_{j=i+1}^{k-1}(1-\beta_{j})\right)^{2}\mathbb{E}\left[\|M^{\prime}_{i+1}\|^{2}\right]$
@@
Here, inequality (a) follows from Assumption 2 and Lemma 3.3, which together imply that $M^{\prime}_{i+1}$ is bounded in expectation by some constant $D_{4}$. Inequality (b) follows from Corollary 2.1.2 from [10] and our assumption that $K_{1}\geq\left(\frac{2\mathfrak{b}}{\beta}\right)^{1/1-\mathfrak{b}}$.
@@
$\sum_{j=i}^{k-1}\beta_{j}\alpha_{j}$ $\leq\sum_{j=i}^{\infty}\frac{\alpha\beta}{(j+K_{1})^{\mathfrak{a}+\mathfrak{b}}}\leq\sum_{j=i}^{\infty}\frac{\alpha\beta}{(j+1)^{\mathfrak{a}+\mathfrak{b}}}\leq\alpha\beta\left(\frac{1}{(i+1)^{\mathfrak{a}+\mathfrak{b}}}+\int_{x=i+1}^{\infty}\frac{1}{x^{\mathfrak{a}+\mathfrak{b}}}dx\right)$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\alpha\beta\left(\frac{1}{(i+1)^{\mathfrak{a}+\mathfrak{b}}}+\frac{1}{(\mathfrak{a}+\mathfrak{b}-1)(i+1)^{\mathfrak{a}+\mathfrak{b}-1}}\right)\stackrel{{\scriptstyle(b)}}{{\leq}}\frac{2\alpha\beta}{\mathfrak{a}+\mathfrak{b}-1}\frac{1}{(i+1)^{\mathfrak{a}+\mathfrak{b}-1}}.$
Here, inequality (a) and (b) both follow from the fact that $1<\mathfrak{a}+\mathfrak{b}<2$, which follows from our assumption that $\alpha_{k}$ and $\beta_{k}$ are both non-summable but square summable sequences. Next,
$\sum_{i=0}^{k}\beta_{i}\sum_{j=i}^{k-1}\beta_{j}\alpha_{j}\leq\frac{2\alpha\beta^{2}}{\mathfrak{a}+\mathfrak{b}-1}\sum_{i=0}^{k-1}\frac{1}{(i+1)^{\mathfrak{a}+2\mathfrak{b}-1}}\stackrel{{\scriptstyle(c)}}{{\leq}}\frac{2\alpha\beta^{2}}{\mathfrak{a}+\mathfrak{b}-1}\sum_{i=0}^{\infty}\frac{1}{(i+1)^{4\mathfrak{a}-1}}\stackrel{{\scriptstyle(d)}}{{\eqqcolon}}D_{5}.$
Inequality (c) follows from the fact that $2\mathfrak{b}\geq 3\mathfrak{a}$, which follows from our assumption that $\beta_{k}^{2}\leq\alpha_{k}^{3}$. Finally, inequality (d) follows from the fact that $4\mathfrak{a}-1>1$, which follows from our assumption that $\mathfrak{a}>0.5$ or $\alpha_{k}$ is square summable.
@@
where $C_{2}\coloneqq(2\Gamma_{6}+2\Gamma_{7}D_{5})/\beta+8\Gamma_{5}\beta$. Here, the first inequality follows from (A.5), and the last inequality follows from the fact that $1/(k+1)^{\mathfrak{b}}\leq 1/(k+1)^{1-\mathfrak{b}}$, which follows from our assumption that $\mathfrak{b}>1/2$. This completes the proof.